What do we do with F=ma when the mass, m, is actually a moving fluid (i.e. air)? Hint: This is used to derive the fundamental equation for Thrust.
Answer #1

from Newton's second law

F=ma

F=dP/dt

F=Pf-Pi/dt

considering both the parts of masses \bigtriangleupM ans M-\bigtriangleupM as forming one and the same system we can write

F=Pf-Pi/dt={(M+\bigtriangleupM)(V+\bigtriangleupV)+\bigtriangleupMu}-[Mu]/dt

where pf and pi are the final and initial momentum of the system respectively

F=M\bigtriangleupv/\bigtriangleupt-V \bigtriangleupM/\bigtriangleupt-\bigtriangleupV *\bigtriangleupM/\bigtriangleupt+u *\bigtriangleupM/\bigtriangleupt

here we observed

\bigtriangleupV*\bigtriangleupM/\bigtriangleupt is not possible

\bigtriangleupM/\bigtriangleupt=-dM/dt

M*\bigtriangleupV/\bigtriangleupt=0

the force F=Mdv/dt+v dM/dt-udM/dt

=0+vdM/dt-u dM.dt

the thrust force =dM/dt(V-u)